OpenStudy (wolfmegan2612):

Solve using mental Math 4y + 3=19 a. 4 b.5 c.5.5 d.64

OpenStudy (anonymous):

Move the 3 to the other side of the equal sign 4y=19-3 4y=16 Divide both sides by 4 4y/4=16/4 y=4

OpenStudy (wolfmegan2612):

Thanks

OpenStudy (anonymous):

@EducateMe there you go haha...

OpenStudy (anonymous):

Gosh, It's like everyone is doing it. Thx